How do I solve this?

When the temperature of 2.35 mA3 of a liquid is increased by 48.5°C it expands by 0.0920 mA3. What is its coefficient of volume expansion? (The answer is 10A-4 CA-1 Just fill in the blank.)
Answer #1

Solution) Vo = 2.35m^3

dV = 0.0920m^3

dT = 48.5°C

We have

dV = (coefficient of volume expansion)(Vo)(dT)

Coefficient of volume expansion = (dV)/(VodT)

Coefficient of volume expansion = (0.0920)/(2.35×48.5)=8.07×10^(-4) C^(-1)